A particle of mass. M moves on straight line with its velocity varying with the distance travelled according to the equation v=a√x where alpha is constant.find the total work done by all the forces during a displacement from x=0 to x=d

Grade:12th pass

1 Answers

Vikas TU
7 years ago
dW = Fdx
dW = Madx 
dW =  Mvdv/dx  * dx
  dW = M * a√x * a/2√x dx
dW = M * a^2/2 dx
Integerating both sides with limits x =0 to x = d we get,
W = (Ma^2/2)d – 0
     = (Ma^2/2)d Joule.